OpenStudy (savannah_noelle):

Graph the line y+5=1/3(x-4) I'll give a medal

OpenStudy (savannah_noelle):

@ParthKohli @pooja195 @Kainui

OpenStudy (denonakavro):

first simplify y+5=1/3(x-4) y+5=1/3x-1.33 y=1/3x-6.3333333 Now graph this in desmos

OpenStudy (savannah_noelle):

How do I graph that?

OpenStudy (denonakavro):

Go here https://www.desmos.com/calculator

OpenStudy (savannah_noelle):

so for the first ordered pair is it (0,-6.3333333) and then the slope is just 1/3 from there?

OpenStudy (denonakavro):

First tell me did you graph it ?

OpenStudy (savannah_noelle):

Yes, is that correct?

OpenStudy (savannah_noelle):

I graphed the point (0,-6.3333333) and made the slope 1/3

OpenStudy (denonakavro):

Then good job

OpenStudy (savannah_noelle):

Oh okay, thank you :)